OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (anonymous):

I got y = -1/2x - 2, is that correct?

OpenStudy (ghazi):

let me check

OpenStudy (anonymous):

slope is incorrect

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

is the slope -5?

OpenStudy (anonymous):

i did the negative reciprical correctly

OpenStudy (ghazi):

slope would be -1/2

OpenStudy (anonymous):

slope is -2 for the equation so it has to be 1/2

OpenStudy (anonymous):

equation is y = –2x + 4

OpenStudy (anonymous):

negative of -2 would be 2 then recipocal of 2 is 1/2

OpenStudy (ghazi):

y-4= 1/2(x+2) solve this

OpenStudy (anonymous):

y-4= 1/2x + 1

OpenStudy (anonymous):

y= 1/2x +5

OpenStudy (ghazi):

nope

OpenStudy (anonymous):

where did i mess up??

OpenStudy (anonymous):

... looks fine to me

OpenStudy (ghazi):

well it's y-4 = 0.5*(x+2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

same thing...

OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4? So would the answer be y= 1/2x + 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

no sorry it's 2y-8=x+2

OpenStudy (ghazi):

2y= x+10

OpenStudy (anonymous):

divide the 2 so you get slope intercept form y = (1/2)x +5